OpenStudy (anonymous):

factoring by groupings: 1.2x raise to 2-x +2xy-y 2.15xy+20y-12x-16 can you help me .

OpenStudy (anonymous):

Is 1. \[2x^2-x+2xy-y\] Please confirm then I can help :-)

OpenStudy (anonymous):

the ans. is: (3x+4)(5y-4)

OpenStudy (anonymous):

@marryrizabolofer I don't think that is the answer....can you show your work.

OpenStudy (callisto):

For question 1, \[2x^2-x +2xy-y\]\[=x(2x-1) +y(2x-1)\]\[=(x+y)(2x-1)\] For question 2, 15xy+20y-12x-16 = 5y ( 3x + 4) - 4 ( 3x+4) = ...? Just take out the common factor and group the rest of the terms, like what I've done for question 1.